OpenStudy (anonymous):

find f'x when f(x)= (2x^7-x^2)(x-1/x+1)

zepdrix (zepdrix):

:o

zepdrix (zepdrix):

\[\Large\bf f(x)=(2x^7-x^2)\left(\frac{x-1}{x+1}\right)\] \[\Large\bf f'(x)=\color{#DD4747}{\left(2x^7-x^2\right)'}\left(\frac{x-1}{x+1}\right)+(2x^7-x^2)\color{#DD4747}{\left(\frac{x-1}{x+1}\right)'}\]

zepdrix (zepdrix):

So we start with product rule

OpenStudy (anonymous):

okay i did that lol

zepdrix (zepdrix):

So we have to take the derivative of the red parts. Can you do the first one?

OpenStudy (anonymous):

yes i can , (14x^7-2x) and (2x/(x+1)^2)?

zepdrix (zepdrix):

14x^6-2x

zepdrix (zepdrix):

Applying Quotient rule: \[\Large\bf \color{#DD4747}{\left(\frac{x-1}{x+1}\right)'}\quad=\quad \frac{(1)(x+1)-(x-1)(1)}{(x+1)^2}\] Hmm I think your numerator is off, our x's will cancel out, yes?

zepdrix (zepdrix):

\[\Large \bf =\frac{2}{(x+1)^2}\]

OpenStudy (anonymous):

oh okay yeah it was lol

OpenStudy (anonymous):

so after that do we find a common denominator?

zepdrix (zepdrix):

So we're here?\[\large\bf f'(x)=\color{royalblue}{\left(14x^6-2x\right)}\left(\frac{x-1}{x+1}\right)+(2x^7-x^2)\color{royalblue}{\left(\frac{2}{(x+1)^2}\right)}\]

zepdrix (zepdrix):

Mmmm yah I guess you could get a common denominator if you want. It's best to just stop here though. If your teacher wants it simplified further though, yes we would want a common denominator of (x+1)^2. So the first term needs another factor of (x+1) in the top and bottom.

OpenStudy (anonymous):

i think ill stop here because itll be too much work lol

zepdrix (zepdrix):

XD good idea haha